OpenStudy (joftheworld):

The company also discovered that it costs $15.50 to produce 3 widgets, $23.50 to produce 7 widgets, and $56 to produce 12 widgets. What is the total cost producing 5 widgets?

OpenStudy (anonymous):

It's a quadratic function, and you have three given points. The form of a quadratic function is ax2+bx+c=0 You have three given (x, y) points: (3, 15.50) (7, 23.5) (12, 56) Insert each of the three points in the equation. You will get three equations in three unknowns, a, b, and c. Solve the system of equations for a, b, and c. Then insert a, b, and c in the quadratic equation above. Finally, evaluate it for x = 5.

OpenStudy (joftheworld):

yeah i know those points. but i dont know which variable to plug them into.

OpenStudy (joftheworld):

@wrstlr3232

OpenStudy (joftheworld):

@SolomonZelman

OpenStudy (anonymous):

\[9a+3b+c = 15.5\] \[49a+7b+c=23.5\] \[144a+12b+c=56\] solve the system

OpenStudy (joftheworld):

a= 1/2 b= -3 c= 20 ????

OpenStudy (anonymous):

a and b are right. I got 19.5 for c

OpenStudy (joftheworld):

how ?

OpenStudy (joftheworld):

its 20.. i just calculated it

OpenStudy (anonymous):

Now substitute 5 into the quadratic equation: 1/2(25)+-3(5)+20 = ? The cost to produce 5 widgits

OpenStudy (joftheworld):

i dont get that part..

OpenStudy (anonymous):

\[ax ^{2}+bx+c=0\]

OpenStudy (anonymous):

1/2(5^2)+-3(5)+20=0 12.5-15+20 = 17.50 so it costs 17.50 to produce 5 widgits (5, 17.50)

OpenStudy (anonymous):

a=1/2 b=-3 c=20 substituted those values for a, b, c...used 5 for the x value since that's what we're looking for

OpenStudy (joftheworld):

thanks can you help me with one more problem?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Looks very similar to last problem

OpenStudy (joftheworld):

any idea ?

OpenStudy (anonymous):

4a+2b+c=45 16a+4b+c=143 100a+10b+c=869 solve the systems

OpenStudy (anonymous):

Let me know what you got for a, b, & c...I've determined those

OpenStudy (anonymous):

Is there any more to the question?\[c(x)=9x ^{2}-5b+19\]

OpenStudy (joftheworld):

a= 9 b= -5 c= 19

OpenStudy (anonymous):

Yes...Great.

OpenStudy (joftheworld):

yeah i have to find the total cost of producing 1 calculator

OpenStudy (anonymous):

substitute a 1 in for x in the c(x) equation I showed up a few

OpenStudy (anonymous):

c(x)=9x2−5b+19 c(x) = 9(1)-5(1)+19 =

OpenStudy (joftheworld):

23?

OpenStudy (anonymous):

yes